a. Using the pairs of values for all 10 ​points, find the equation of the regression line.

b. After removing the point with coordinates ​(7,1), use the pairs of values for the remaining 9 points and find the equation of the regression line.

c. Compare the results from parts​ (a) and​ (b).

a. What is the equation of the regression line for all 10 ​points?

y=___+___x​ (Round to three decimal places as​ needed.)

b. What is the equation of the regression line for the set of 9 points?

y=___ ​(Round to three decimal places as​ needed.)

c. Choose the correct description of the results below.

A. The regression line is very similar in both cases.

B. The removal of the point has a significant impact on the regression line.

C. The regression line​ changes, but the change is small.

D. There is no regression line for the second case because the data are in a pattern.

ф ку 10- 84 6- 4- 24 х 0 0 2 4 6 8 10
Answer #1

n ac x2 y 2 ny Cur (1,81 (418) - uq 7 L 8 by (217) 8 9 - 81 a 2 7 4 49 14 2 (218) (2,4) (317) (3,82 (3,9) (711) 4 64 16 4 81Q.2 y-1.1052+ 10.063 Answer. b.) New values, Ex=25-7= 18 and n=a. =73-1 =72 zy z = = 2, 9 = 72 = 8. que 9+49 = 42 2y? = 583-